OpenStudy (anonymous):

1(6x^2 - 11x -7) can you check for GCF and factor completely ????

OpenStudy (saifoo.khan):

GCF = 1

OpenStudy (saifoo.khan):

\[(2 x+1) (3 x-7) \]

OpenStudy (anonymous):

there is no GCF 6x^2 - 11x -7 =6x^2+3x-14x-7 =3x(2x+1)-7(2x+1) =(2x+1)(3x-7)

OpenStudy (anonymous):

saifoo is correct GCF is 1

OpenStudy (anonymous):

got it thanks..